OpenStudy (anonymous):

Multi-Step equation 3(x + 2) = 6 Need help solving 3x+6=6 3x-6-6 both of the 6's cancel each other out,but what do I do with the 3 now?

OpenStudy (anonymous):

So you have 3x = 0 Divide both sides by 3, you get x = 0/3 = 0 x = 0.

OpenStudy (anonymous):

3(x + 2) = 6 You multiply 3 by x and 2 3x+6=6 -6 -6 3x=0 /3 /3 x=0

OpenStudy (anonymous):

Had a feeling it was 0 but I was unsure thanks for the help

OpenStudy (anonymous):

Welcome.